SUPERSONIC FREE-JET QUANTUM CASCADE LASER MEASUREMENTS OF ν_{4} FOR CF_{3}^{35}Cl AND CF_{3}^{37}Cl AND FTS MEASUREMENTS FROM 400 TO 1260 cm^{−1}

JAMES F. KELLY, \underline{THOMAS A. BLAKE}, ROBERT L. SAMS, Pacific Northwest National Laboratory, P. O. Box 999, Mail Stop K8-88, Richland, WA 99352 (PNNL is operated for the US Department of Energy by the Battelle Memorial Institute under contract DE-AC05-76RL0 1830); ARTHUR MAKI, 15012 24th Ave. S.E., Mill Creek, WA 98012.

A supersonic free-jet spectrum of the ν_{4} band of CF_{3}Cl has been measured using a quantum cascade laser system. Those measurements were combined with a low temperature (-67 C) FTS spectrum of the region 1060 to 1260 cm^{−1} to give improved values for the rovibrational constants for the ν_{1}, 2ν_{5}, and ν_{4} states of the CF_{3}^{35}Cl AND CF_{3}^{37}Cl. The principal perturbation found by earlier investigators in the ν_{1} band is treated as a very weak Coriolis interaction at several avoided crossings of the rotational levels of the ν_{1} state and the 2ν_{5} state with kl < 0. Room temperature FTS measurements were also made for the region 400 to 970 cm^{−1}. With these new measurements we now have high resolution data on the states ν_{1}, ν_{2}, ν_{3}, ν_{4}, and ν_{5}. Also included in this analysis are the overtones 2ν_{3} and 2ν_{6} and a few hot bands.